NEW ABSORPTION SPECTRA OF CH_{2} NEAR 780 NM

JU XIN, Department of Physics and Engineering Technology, Bloomsburg University, 400 East Second Street, Bloomsburg PA 17815-1301; \underline{ZHONG WANG} and TREVOR J. SEARS, Department of Chemistry, Brookhaven National Laboratory, Upton, New York 11973 and Department of Chemistry, Stony Brook University, Stony Brook, New York 11794.

The near infrared and visible spectrum (b̃^{1}B_{1} − ã^{1}A_{1}) of singlet CH_{2} has been the subject of much study. However, the region between the red end of the visible part of the spectrum and about 800 nm has not been recorded since the pioneering work of Herzberg and Johns.^{a} We have remeasured the absorption spectrum between approximately 769 and 806 nm at near shot-noise-limited sensitivity and Doppler-limited resolution using a frequency-modulated extended cavity diode laser source. Rotational branches in 7 vibronic bands involving K_{a} = 0 − 4 have been assigned using known ground state combination differences. Most of them have not previously been observed and some reassignments of the Herzberg and Johns analysis have been made. Comparison with the most complete available calculated ro-vibronic energy level structure^{b} helped considerably in making the assignments, and the observed vibronic levels are assigned to levels of both ã and b̃ electronic character. The calculated energy levels show moderate, up to 10 cm^{−1}, apparently random, differences from the observed levels The new data will certainly help to refine the singlet potential and also provide additional avenues for future kinetics and dynamics studies of the radical. Acknowledgments: Work at Brookhaven National Laboratory was carried out under Contract No. DE-AC02-98CH10886 with the U.S. Department of Energy. Ju Xin acknowledges support from the Faculty and Student Teams program of the Educational Programs Department at Brookhaven National Laboratory.

aG. Herzberg and J. W. C. Johns Proc. R. Soc. London Ser. A, 295, 107 (1966)

bJ. -P. Gu, G. Hirsch, R J Buenker, M. Brumm, G. Osmann, P. R. Bunker and P. Jensen J. Molec. Struc., 517-8, 247 (2000)